OpenStudy (anonymous):

Can someone PLEASE explain to me how I got this wrong and can you explain the correct answer? I'd really appreciate it. 12. What is the y-intercept (or 'b') for the equation that contains the coordinates (3,3) and (2,-2)? A. -12 B. 5 C. 12 D. 1 I chose 5 for the answer, and it is INCORRECT.

OpenStudy (anonymous):

Did you find your equation using point-slope formula?

OpenStudy (anonymous):

I can help you with this one.

OpenStudy (anonymous):

Remember, that the Y intercept is where a given point in a function intersects on the Y-axis on a plane. So. We first need an equation. -2 - 3 = -5 2 - 3 = -1 So your slope is 5. Pick the first point. y - 3 = 5(x-3) There's your equation. Now, plug in a Y-intercept to find if it solves the equation. 0,1 1-3 = 5 (0-3) is -2 = -15? No. 0,12 12- 3 = 5 ( 0-3 ) is 9 = -15? No. 0,-12 -12 - 3 = 5 ( 0-3 ) Is -15 = -15? Yes, So your answer is (0,-12).

OpenStudy (anonymous):

Bye, have a nice day.